Mix a solution that is 40% alcohol with a solution that is 10% alcohol to make 225 mL of a solution that is 20% alcohol. How much of each solution should you use?

Answer by rwm(914)

x+y=225
x=225-y
.4x+.1y=225*.2
.4(225-y)+.1y=225*.2
solve for y
then for x